# FastDraft: How to Train Your Draft

### v2

@@ 1 Introduction
The advent of Transformer architectures has fundamentally reshaped the field of natural language processing (NLP). In recent years, Transformer-based models have achieved remarkable success across a broad spectrum of natural language understanding and generation tasks (Achiam et al., 2023; Team et al., 2023). Their exceptional performance, particularly in large language models (LLMs), has made them highly desirable for deployment in numerous applications, ranging from conversational systems to content generation and beyond. Despite their outstanding performance, LLMs suffer from slow inference speed due to substantial memory bandwidth requirements and the sequential nature of auto-regressive generation (ARG). The introduction of Speculative Decoding (SD) (Leviathan et al., 2023) offers a promising solution for accelerating ARG without sacrificing generation quality, making it a compelling approach for improving LLM inference efficiency. SD utilizes a draft language model (LM) to generate a sequence of tokens auto-regressively, while the target model validates the batched tokens in parallel. In certain applications, SD can achieve a 2-3x speedup in LLM inference without compromising the generation quality of the target model. Achieving significant speedup with SD requires a high-quality draft model that is both efficient and well-aligned with the target. To date, such draft models remain scarce, even for widely used open-source LLMs (Dubey et al., 2024; Abdin et al., 2024) due to vocabulary incompatibility.
To address this limitation, we propose FastDraft, a method for producing hardware (HW) efficient draft models that are orders of magnitude smaller than their corresponding target models.
@@ 2 Related work
The widespread adoption of LLMs in cloud and edge devices has driven a significant body of research focused on developing alternative strategies for ARG to address the slow performance of LLM inference (Santilli et al., 2023; Ghazvininejad et al., 2019; Stern et al., 2018). However, many of these approaches compromise generation quality or require additional training data and architectural modifications. The introduction of speculative decoding as a lossless solution for accelerating LLM inference (Leviathan et al., 2023) has inspired a new wave of follow-up research. Some studies propose using plug-in prediction heads as a drafting mechanism (Zhang et al., 2024; Cai et al., 2024), while others focus on improving the serving latency of stand-alone draft models (Sun et al., 2024; Chen et al., 2023; Miao et al., 2023). In contrast, our work focuses on directly enhancing stand-alone draft model’s capabilities through pre-training and fine-tuning. Other works apply KD to draft models to improve alignment with the target model. These studies explore various divergence functions for the KD algorithm, rather than relying solely on the commonly used Kullback–Leibler Divergence (KLD). For instance, Zhou et al. (2023) proposed Total Variation Distance (TVD) based on Corollary 3.6 in (Leviathan et al., 2023), which posits that minimizing TVD maximizes the token-level acceptance rate. Goel et al. (2024) further developed this approach with TVD++, fine-tuning their model, pre-trained on 600 billion tokens. Another study, Yan et al. (2024), focuses on enhancing the HW efficiency of draft models by extensively analyzing the trade-off between time latency and acceptance rates, rather than relying solely on the latter. While their work emphasizes optimizing the draft architecture, our approach fixes the architecture and introduces efficient approach for pre-training and aligning a draft model. Li et al. (2024) proposes training a compact draft model based on the target model, using a relatively limited dataset. The primary aim of the paper is to develop a draft architecture that effectively utilizes the target model’s hidden representations and weights. In contrast, our paper centers on training and aligning any draft architecture that shares only the vocabulary with the target model.
@@ 5 Experiments
In this section we report descriptions, setup and results of our experiments. In the following experiments, we report acceptance rate (Section 4.3.1) results using two sampling methods, greedy and multinomial sampling decoding with temperature $T=0.6$ with block sizes $\gamma=3$ and $\gamma=5$ unless stated otherwise. The hyper-parameters used in our draft pre-training experiments are detailed Appendix A.3 unless stated otherwise.
@@ 5.1 Experimental setup
For our natural language dataset, we use a 10 billion tokens (BT) sample from the FineWeb dataset (Penedo et al., 2024). FineWeb is a widely used open-source, de-duplicated, and quality-filtered dataset, comprising 15 trillion tokens derived from 96 Common Crawl snapshots (Crawl, ). It has been demonstrated to yield better-performing LLMs compared to other open-source pre-training datasets such as C4, RedPajama, and The Pile (Section 3.7 in Penedo et al. (2024)). For our code dataset, we use a 10BT sample from The Stack v2 smol dataset (Lozhkov et al., 2024). This variant of The Stack v2 consists of 17 commonly used programming languages, as well as a substantial collection of documentation languages, configuration languages, and configuration files. An overview of the 10BT dataset composition is provided in Appendix A.2.
@@ 5.3 Continued pre-training for code
Given the structured nature of code, it serves as a highly suitable domain for generation with SD. To investigate how to optimize pre-training for code drafts, we conducted a comprehensive ablation study. This study involved pre-training from scratch on a mixture of code and pre-training with CP approach, where the base draft model was pre-trained on code, and CP was applied to adapt the model for natural language instead of code. Detailed results and analysis of this ablation study are provided in Appendix A.5. Overall, our findings indicate that initializing continued pre-training from a text-based model using a mixed CP dataset offers the most effective pre-training strategy for draft models, compared to other configurations tested.
@@ 5.6 Hardware-aware draft
The Phi3-mini 50M uses the original Transformers dimensions, however, due to the strict latency requirements of draft models, their architectural design warrants special attention, as their performance may vary depending on the hardware platform. We conduct a series of experiments to examine the performance of draft models derived from different architectural choices. Our focus is on altering the depth of the models by changing the number of layers or the width of the models by modifying the hidden dimension. The target system for our experiments is the latest Intel® Core Ultra series 200 (Lunar Lake) processors coupled with Intel® OpenVINO runtime engine.
First, we analyze the effect of scaling the width and depth of the draft on latency of the draft model on the target hardware. We fix either the hidden dimension or the model depth to match the corresponding values pf Phi3-mini 50M draft and modify the other parameter to observe how latency scales with wider or deeper models. As shown in Figure 3, our benchmarking results suggest that increasing the number of layers has the most significant impact on latency, while increasing the hidden dimension leads to only a slight increase.
Next, we evaluate the impact of the draft architecture under a fixed parameter budget4 4 4 We excluded embedding table size from the model parameter budget, as it does not strain memory bandwidth. by assessing the AR and estimated speedup achieved when varying the width and depth of the draft. By redistributing parameters between these dimensions, we aimed to identify configurations that optimize the trade-off between speculation quality and computational efficiency. We pre-trained multiple draft models for Phi-3-mini, varying the number of layers and hidden dimension sizes while maintaining a fixed parameter count. We then measured their AR on a target dataset and recorded their latency on our target hardware to estimate the speedup achieved through SD via Equation 2. Figure 4 presents the results for the CNN-DM dataset. Our findings indicate that increasing depth does not improve the AR, while excessively shallow drafts degrade performance. Moreover, the AR–latency trade-off follows an inverted U-shaped speedup curve with a broad base, consistent with our earlier experimental results. The results for the TS and Dolly datasets, presented in Appendix B, exhibited the same pattern.
@@
Figure 4: Estimated speedup on CNN-DM for models with constant parameter budget. L, H denotes the number of layers and hidden dimension size
@@ 6 Results & reproducibility
Combining the findings from the ablation studies presented in Section 5, we outline our comprehensive pipeline for draft pre-training and fine-tuning to optimize performance on key metrics for speculative decoding and demonstrate it by producing a draft for Phi-3-mini. To further illustrate reproducibility, we employ the same pipeline to produce a draft for Llama-3.1-8B-Instruct Dubey et al. (2024). We present a 50M draft for the Phi-3-mini and a 150M draft for Llama-3.1-8B. Drafts’ architecture details are presented in Appendix A.1. Considering both performance and resource efficiency, our findings from sections 5.2, 5.3 suggest pre-training over 5BT of FineWeb text data and continued pre-training on a mixture of 5BT code data from The Stack v2 and 2.5BT FineWeb text data as a best practice. For fine-tuning, we conclude from sections 5.4, 5.5 that sequence-level KD yields significant improvements while token-level KD benefits are not definitive, therefore, for FastDraft we only utilize sequence-level KD. We construct an alignment dataset for both Phi-3 and Llama-3.1 drafts by combining a number of synthetic datasets we generated with the appropriate target model as described in Section 5.1. Table 4 presents the AR improvements achieved through each stage of FastDraft generated with multinomial sampling. Results for greedy sampling are presented in Table 17 in the Appendix. Using both sampling methods, we observe for datasets CNN-DM, TS, and Dolly, a substantial AR increase following fine tuning, with instruction-following dataset Dolly exhibiting an increase of ~10% in AR. For HumanEval the primary performance gains stem from pre-training on code domain knowledge during continued pre-training. These models achieve a MBSU of ~2x for natural language tasks and ~3x for code completion tasks.
Furthermore, we demonstrate the real-world wall-clock time speedup achieved with FastDraft on the Intel® Core Ultra series 200 as detailed in Section 5.6. We observe an average TPOT speedup of 1.5x for summarization tasks and 2x for code completion. Additionally, memory transaction efficiency improves, with up to 3x reduction in memory bandwidth usage, as indicated by the MBSU metric in these trials. Memory transactions can be highly energy-intensive (Horowitz, 2014), making the use of FastDraft drafts a promising approach for achieving substantial power savings on edge devices. These performance gains stem from the compact nature of our drafts, enabling them to operate in tandem with large target LLMs. We believe FastDraft has the potential to unlock LLMs inference on AI-PC and other edge-devices.
@@ 8 Limitations
Appendix A Pre-training details and results
A.1 Draft model architecture
@@
A.2 Details of pre-training code dataset
@@
A.3 Hyper-parameter configuration for datasets size in pre-training experiments
@@
A.4 Pre-training model size ablation
@@
A.5 Pre-training on code ablation
@@
Appendix B Hardware-aware draft details and results
@@
constant parameter budget. L, H denotes the number
@@
constant parameter budget. L, H denotes the number
@@
Appendix C FastDraft additional results

### v3

@@ 1 Introduction
The advent of Transformer architectures has fundamentally reshaped the field of natural language processing (NLP). In recent years, Transformer-based models have achieved remarkable success across a broad spectrum of natural language understanding and generation tasks (Achiam et al., 2023; Team et al., 2023). Their exceptional performance, particularly in large language models (LLMs), has made them highly desirable for deployment in numerous applications, ranging from conversational systems to content generation and beyond. Despite their outstanding performance, LLMs suffer from slow inference speed due to substantial memory bandwidth requirements and the sequential nature of auto-regressive generation (ARG). The introduction of Speculative Decoding (SD) (Leviathan et al., 2023) offers a promising solution for accelerating ARG without sacrificing generation quality, making it a compelling approach for improving LLM inference efficiency. SD utilizes a draft language model (LM) to generate a sequence of tokens auto-regressively, while the target model validates the batched tokens in parallel. In certain applications, SD can achieve a 2-3x speedup in LLM inference without compromising the generation quality of the target model. Achieving significant speedup with SD requires a high-quality draft model that is both efficient and well-aligned with the target. To date, such draft models remain scarce, even for widely used open-source LLMs (Dubey et al., 2024; Abdin et al., 2024) due to vocabulary incompatibility. To address this limitation, we propose FastDraft, a method for producing hardware (HW) efficient draft models that are orders of magnitude smaller than their corresponding target models.
@@ 2 Related work
The widespread adoption of LLMs in cloud and edge devices has driven a significant body of research focused on developing alternative strategies for ARG to address the low inference efficiency of LLMs (Santilli et al., 2023; Ghazvininejad et al., 2019; Stern et al., 2018). However, many of these approaches compromise generation quality or require additional training data and architectural modifications. The introduction of speculative decoding as a lossless approach for accelerating LLM inference (Leviathan et al., 2023) has inspired a new wave of follow-up research. Some studies propose using plug-in prediction heads as a drafting mechanism (Zhang et al., 2024; Cai et al., 2024), while others focus on improving the serving latency of stand-alone draft models (Sun et al., 2024; Chen et al., 2023; Miao et al., 2023). In contrast, our work focuses on directly enhancing stand-alone draft model’s capabilities through pre-training and fine-tuning. We provide more details on the comparison with these methods in Appendix A. Other works apply KD to draft models to improve alignment with the target model. These studies explore various divergence functions for the KD algorithm, rather than relying solely on the commonly used Kullback–Leibler Divergence (KLD). For instance, Zhou et al. (2023) proposed Total Variation Distance (TVD) based on Corollary 3.6 in (Leviathan et al., 2023), which posits that minimizing TVD maximizes the token-level acceptance rate. Goel et al. (2024) further developed this approach with TVD++, fine-tuning their model, pre-trained on 600 billion tokens. Another study, Yan et al. (2024), focuses on enhancing the HW efficiency of draft models by extensively analyzing the trade-off between time latency and acceptance rates, rather than relying solely on the latter. While their work emphasizes the approach for obtaining a draft model by pruning of the shelf models, our method introduces an efficient approach for pre-training and aligning a draft model. Li et al. (2024) proposes training a compact draft model based on the target model, using a relatively limited dataset. The primary aim of the paper is to develop a draft architecture that effectively utilizes the target model’s hidden representations and weights. In contrast, our paper centers on training and aligning any draft architecture that shares only the vocabulary with the target model.
@@ 5 Experiments
In this section we report descriptions, setup and results of our experiments. In the following experiments, we report acceptance rate (Section 4.3.1) results using two sampling methods, greedy and multinomial sampling decoding with temperature $T=0.6$ with block sizes $\gamma=3$ and $\gamma=5$ unless stated otherwise. The hyper-parameters used in our draft pre-training experiments are detailed Appendix B.3 unless stated otherwise.
@@ 5.1 Experimental setup
For our natural language dataset, we use a 10 billion tokens (BT) sample from the FineWeb dataset (Penedo et al., 2024). FineWeb is a widely used open-source, de-duplicated, and quality-filtered dataset, comprising 15 trillion tokens derived from 96 Common Crawl snapshots (Crawl, ). It has been demonstrated to yield better-performing LLMs compared to other open-source pre-training datasets such as C4, RedPajama, and The Pile (Section 3.7 in Penedo et al. (2024)). For our code dataset, we use a 10BT sample from The Stack v2 smol dataset (Lozhkov et al., 2024). This variant of The Stack v2 consists of 17 commonly used programming languages, as well as a substantial collection of documentation languages, configuration languages, and configuration files. An overview of the 10BT dataset composition is provided in Appendix B.2.
@@ 5.3 Continued pre-training for code
Given the structured nature of code, it serves as a highly suitable domain for generation with SD. To investigate how to optimize pre-training for code drafts, we conducted a comprehensive ablation study. This study involved pre-training from scratch on a mixture of code and pre-training with CP approach, where the base draft model was pre-trained on code, and CP was applied to adapt the model for natural language instead of code. Detailed results and analysis of this ablation study are provided in Appendix B.5. Overall, our findings indicate that initializing continued pre-training from a text-based model using a mixed CP dataset offers the most effective pre-training strategy for draft models, compared to other configurations tested.
@@ 5.6 Hardware-aware draft
The Phi3-mini 50M uses the original Transformers dimensions, however, due to the strict latency requirements of draft models, their architectural design warrants special attention, as their performance may vary depending on the hardware platform. We conduct a series of experiments to examine the performance of draft models derived from different architectural choices. Our focus is on altering the depth of the models by changing the number of layers or the width of the models by modifying the hidden dimension. The target system for our experiments is the latest Intel® Core Ultra 7 Processor 268V coupled with Intel® OpenVINO runtime engine.
First, we analyze the effect of scaling the width and depth of the draft on latency of the draft model on the target hardware. We fix either the hidden dimension or the model depth to match the corresponding values of Phi3-mini 50M draft and modify the other parameter to observe how latency scales with wider or deeper models. As shown in Figure 3, our benchmarking results suggest that increasing the number of layers has the most significant impact on latency, while increasing the hidden dimension leads to only a slight increase.
Next, we evaluate the impact of the draft architecture under a fixed parameter budget4 4 4 We excluded embedding table size from the model parameter budget, as it does not strain memory bandwidth. by assessing the AR and estimated speedup achieved when varying the width and depth of the draft. By redistributing parameters between these dimensions, we aimed to identify configurations that optimize the trade-off between speculation quality and computational efficiency. We pre-trained multiple draft models for Phi-3-mini, varying the number of layers and hidden dimension sizes while maintaining a fixed parameter count. We then measured their AR on a target dataset and recorded their latency on our target hardware to estimate the speedup achieved through SD via Equation 2. Figure 4 presents the results for the CNN-DM dataset. Our findings indicate that increasing depth does not improve the AR, while excessively shallow drafts degrade the performance. Moreover, the AR–latency trade-off follows an inverted U-shaped speedup curve with a broad base, consistent with our earlier experimental results. The results for the TS and Dolly datasets, presented in Appendix C, exhibited the same pattern.
@@
Figure 4: Estimated speedup on CNN-DM for models with constant parameter budget. L, H denote the number of layers and hidden dimension size
@@ 6 Results & reproducibility
Combining the findings from the ablation studies presented in Section 5, we outline our comprehensive pipeline for draft pre-training and fine-tuning to optimize performance on key metrics for speculative decoding and demonstrate it by producing a draft for Phi-3-mini. To further illustrate reproducibility, we employ the same pipeline to produce a draft for Llama-3.1-8B-Instruct Dubey et al. (2024). We present a 50M draft for the Phi-3-mini and a 150M draft for Llama-3.1-8B. Drafts’ architecture details are presented in Appendix B.1. Considering both performance and resource efficiency, our findings from sections 5.2, 5.3 suggest pre-training over 5BT of FineWeb text data and continued pre-training on a mixture of 5BT code data from The Stack v2 and 2.5BT FineWeb text data as a best practice. For fine-tuning, we conclude from sections 5.4, 5.5 that sequence-level KD yields significant improvements while token-level KD benefits are not definitive, therefore, for FastDraft we only utilize sequence-level KD. We construct an alignment dataset for both Phi-3 and Llama-3.1 drafts by combining a number of synthetic datasets we generated with the appropriate target model as described in Section 5.1. Table 4 presents the AR improvements achieved through each stage of FastDraft generated with multinomial sampling. Results for greedy sampling are presented in Table 17 in the Appendix. Using both sampling methods, we observe for datasets CNN-DM, TS, and Dolly, a substantial AR increase following fine tuning, with instruction-following dataset Dolly exhibiting an increase of ~10% in AR. For HumanEval the primary performance gains stem from pre-training on code domain knowledge during continued pre-training. These models achieve a MBSU of ~2x for natural language tasks and ~3x for code completion tasks.
Furthermore, we demonstrate the real-world wall-clock time speedup achieved with FastDraft on the Intel® Core Ultra 7 Processor 268V as detailed in Section 5.6. We observe an average TPOT speedup of 1.5x for summarization tasks and 2x for code completion. Additionally, memory transaction efficiency improves, with up to 3x reduction in memory bandwidth usage, as indicated by the MBSU metric in these trials. Memory transactions can be highly energy-intensive (Horowitz, 2014), making the use of FastDraft drafts a promising approach for achieving substantial power savings on edge devices. These performance gains stem from the compact nature of our drafts, enabling them to operate in tandem with large target LLMs. Notably, a pre-trained FastDraft draft model can be used either as-is or aligned with any compatible model. For instance, our Llama-3.1 draft is readily compatible with over 1,000 models on the HuggingFace Hub5 5 5 https://huggingface.co/, highlighting FastDraft’s scalability. We believe FastDraft has the potential to unlock LLMs inference on AI-PC and other edge-devices.
@@ 8 Limitations
Appendix A Comparison to other methods
To ensure a fair comparison with speculative decoding methods such as MEDUSA (Cai et al., 2024), EAGLE (Li et al., 2024), and self-speculation techniques as presented in (Zhang et al., 2023; Xia et al., 2024), key factors must remain consistent, including hardware, target model, and speculation strategy. In contrast to the aforementioned methods, FastDraft adheres strictly to standard, sequential speculative decoding without employing multi-token windows, multiple drafts, or speculation trees—although such enhancements are complementary to our approach and could be incorporated to achieve additional speed improvements. Moreover, our primary objective is to offer a solution in scenarios where utilizing an existing draft model or a smaller language model (SLM) from the same model family is not feasible (e.g., Phi-3).
As reported in Goel et al., 2024, pretrained draft models have been empirically shown to exhibit significantly better alignment with the target model. Despite the associated cost, our approach enables efficient training from scratch in several key aspects: FastDraft supports architectural flexibility, enabling the construction of compact draft models. In contrast, drafts designed to leverage latent representations of the target model scale with the target’s hidden dimension. For example, EAGLE drafts for 7B, 13B, and 70B models require 0.24B, 0.37B, and 0.99B parameters, respectively, while MEDUSA’s 3–5 decoding heads correspond to 1.5–2.6B trainable parameters for Llama3.1-8B. In the case of Self-Speculation, the large budget of target parameters—approximately 500M for Llama-3.1-8B—used during speculation also limits speedup and results in substantial memory consumption. The FastDraft pipeline can be completed within one day using eight accelarators, representing a substantial improvement over prior draft training methods  (Goel et al., 2024; Miao et al., 2023). For comparison, EAGLE training requires up to 2 days on four GPUs. Furthermore, EAGLE-3 (Li et al., 2025) was trained on approximately eight times more data than EAGLE. Moreover, FastDraft drafts can be used with any compatible target model without additional training or adjustments, whereas fine-tuning can further improve alignment to the target model. Self-Speculation, by contrast, depends on Bayesian optimization—a computationally intensive process that can take several hours per iteration.
Appendix B Pre-training details and results
B.1 Draft model architecture
@@
B.2 Details of pre-training code dataset
@@
B.3 Hyper-parameter configuration for datasets size in pre-training experiments
@@
B.4 Pre-training model size ablation
@@
B.5 Pre-training on code ablation
@@
Appendix C Hardware-aware draft details and results
@@
constant parameter budget. L, H denote the number
@@
constant parameter budget. L, H denote the number
@@
Appendix D FastDraft additional results